Session: Nov 2004 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Sun Nov 7 06:41:56 2004) Minor failures : POIN FS
We lost some scans due to pointing/field system errors. At 21.40UT we started regulary

Onsala (Michael Lindqvist at Sun Nov 7 06:54:07 2004) Success :
No known problems.

Torun (K. Borkowski at Mon Nov 8 06:32:36 2004) Success :
No known problems.

Effelsberg (A. Kraus at Mon Nov 8 08:21:14 2004) Success :

Westerbork (Tony Foley at Mon Nov 8 10:51:41 2004) Minor failures : HA-E
Single dish was used. Please change the coordinates in the vexfile accordingly. Many scans outside Wb hour angle range. Othewise OK

Cambridge (Alastair Gunn at Mon Nov 8 10:56:27 2004) Success :
No known problems. Note that the Cambridge formatter was resynced prior to this experiment.

(Other) (Alastair Gunn at Mon Nov 8 10:57:08 2004) Success :
The Darnhall telescope was used. No known problems.

Noto (P.Cassaro at Tue Nov 9 19:12:31 2004) Minor failures :
Formatter error from 6:30 UT. From 7:30 UT to the end no known problems.